Okay, I guess I did not get it. the little 3 outside the square root of 8y^3
please explain.

Answer by scott8148(6628)

the "square" root symbol is just a root symbol where the 2 is the "default" value ... if there is no number, it is a square root

as you have discovered, any positive integer may be used

in this case, a 3 indicates a "cube" root ... something multiplied by itself 3 times to get the value inside the symbol

since 2*2*2=8 and y*y*y=y^3, the cube root of 8y^3 is 2y